# Bayesian Experimental Design via Contrastive Diffusions

### v1

@@ Abstract
Bayesian Optimal Experimental Design (BOED) is a powerful tool to reduce the cost of running a sequence of experiments. When based on the Expected Information Gain (EIG), design optimization corresponds to the maximization of some intractable expected contrast between prior and posterior distributions. Scaling this maximization to high dimensional and complex settings has been an issue due to BOED inherent computational complexity. In this work, we introduce an expected posterior distribution with cost-effective sampling properties and provide a tractable access to the EIG contrast maximization via a new EIG gradient expression. Diffusion-based samplers are used to compute the dynamics of the expected posterior and ideas from bi-level optimization are leveraged to derive an efficient joint sampling-optimization loop, without resorting to lower bound approximations of the EIG. The resulting efficiency gain allows to extend BOED to the well-tested generative capabilities of diffusion models. By incorporating generative models into the BOED framework, we expand its scope and its use in scenarios that were previously impractical. Numerical experiments and comparison with state-of-the-art methods show the potential of the approach.
@@ 1 Introduction
Designing optimal experiments can be critical in numerous applied contexts where experiments are constrained in terms of resources or more generally costly and limited. In this work, design is assumed to be characterized by some continuous parameters ${\boldsymbol{\xi}}\leavevmode\nobreak\ \in\leavevmode\nobreak\ {\cal E}\subset\mathbb{R}^{d}$, which refers to the experimental part, such as the choice of a measurement location, that can be controlled to optimize the experimental outcome. We consider a Bayesian setting in which the parameters of interest is ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}\subset\mathbb{R}^{m}$ and design is optimized to maximize the information gain on ${\boldsymbol{\theta}}$. Bayesian optimal experimental design (BOED) is not a new topic in statistics, see e.g. Chaloner and Verdinelli, (1995); Sebastiani and Wynn, (2000); Amzal et al., (2006) but has recently gained new interest with the use of machine learning techniques, see Rainforth et al., (2024); Huan et al., (2024) for recent reviews. The most common approach consists of maximizing the so-called expected information gain (EIG), which is a mutual information criterion that accounts for information via the Shannon’s entropy. Let $p({\boldsymbol{\theta}})$ denote a prior probability distribution and $p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ a likelihood defining the observation ${\mathb{y}}\in{\cal Y}$ generating process. The prior is assumed to be independent on ${\boldsymbol{\xi}}$ and $p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ available in closed-form. To our knowledge, all previous BOED approaches also assume that the prior is available in closed-form, a setting that we will refer to as density-based BOED. In this work, we will also consider data-based BOED, a setting where only samples from the prior are available. The EIG, denoted below by $I$, admits several equivalent expressions, see e.g. Foster et al., (2019). It can be written as the expected loss in entropy when accounting for an observation ${\mathb{y}}$ at ${\boldsymbol{\xi}}$ (eq. (1)) or as a mutual information (MI) or expected Kullback-Leibler (KL) divergence (eq. (2)). Denoting $p_{\boldsymbol{\xi}}({\boldsymbol{\theta}},{\mathb{y}})=p({\boldsymbol{\theta}},{\mathb{y}}|{\boldsymbol{\xi}})$ the joint distribution of $({\boldsymbol{\theta}},{\mathb{Y}})$ and using $p({\boldsymbol{\theta}},{\mathb{y}}|{\boldsymbol{\xi}})=p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})p({\mathb{y}}|{\boldsymbol{\xi}})=p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})p({\boldsymbol{\theta}})$, it comes,
@@
In this work, we propose a novel 1-stage approach which leverages a sampling-as-optimization setting (Korba and Salim,, 2022; Marion et al.,, 2024) where sampling is seen as an optimization task over the space of probability distributions. We introduce a new EIG gradient expression (Section 3), which highlights the EIG gradient as a function of both the design and some sampling outcome. This fits into a bi-level optimization framework adapted to BOED in Section 4. So doing, at each step, both an estimation of the optimal design and samples from the current posterior distribution can be provided in a single loop described in Section 5. It results an efficient procedure that can handle both traditional density-based samplers and data-based samplers such as provided by the highly successful diffusion-based generative models. The resulting efficiency gain enables BOED applications at significantly larger scales than previously feasible. Notably, this opens the first access to data-based BOED, beyond density-based approaches. Figure 1 is an illustration on a $28\times 28$ image ${\boldsymbol{\theta}}$ reconstruction problem from $7\times 7$ sub-images centered at locations ${\boldsymbol{\xi}}$ to be selected, details in Section 6. For simpler notation, we first present our approach in the static design case. Adaptation to the sequential case is specified in Section 6 and all numerical examples are in the sequential setting.
@@ 2 Related work
We focus on gradient-based BOED for continuous problems. Applying a first-order method to solve (3) requires computing gradients of the EIG $I$, which are no more tractable than $I$ itself. Gradient-based BOED is generally based on stochastic gradient-type algorithms (see Section 4.3.2. in Huan et al., (2024)). This requires in principle unbiased gradient estimators, although stochastic approximation solutions using biased oracles have also been investigated, see e.g. Demidovich et al., (2023); Liu and Tajbakhsh, (2024). To meet this requirement, most stochastic gradient-based approaches start from an EIG lower bound that yields tractable unbiased gradient estimators. More specifically, EIG lower bounds have usually the advantage to remove the nested expectation issue, see e.g. Foster et al., (2019). In contrast, very few approaches focus on direct EIG gradient estimators. To our knowledge, this is only the case in Goda et al., (2022) and Ao and Li, (2024). Goda et al., (2022) propose an unbiased estimator of the EIG gradient using a randomized version of a multilevel nested Monte Carlo (MLMC) estimator from Rhee and Glynn, (2015). A different estimator is proposed by Ao and Li, (2024), who use MCMC samplers leading to biased estimators, for which the authors show empirically that the bias could be made negligible. In this work, we first show, in Section 3, that their two apparently different solutions actually only differ in the way the intractable posterior distribution is approximated. We then propose a third way to compute EIG gradients that is more computationally efficient and scales better to larger data volumes and sequential design contexts. This new expression makes use of a distribution that we introduce and name the expected posterior distribution. This latter distribution has interesting sampling features that allows us to leverage score-based sampling techniques and connect to the so-called implicit diffusion framework of Marion et al., (2024). Our single loop procedure in Section 5 is inspired by Marion et al., (2024) and other recent developments in bi-level optimization (Yang et al.,, 2021; Dagréou et al.,, 2022; Hong et al.,, 2023). However, these latter settings do not cover doubly intractable objectives such as the EIG, which requires both appropriate gradient estimators and sampling operators, see our Sections 3 and 4. In BOED, efficient single loop procedures have been proposed by Foster et al., (2020) but they rely heavily on variational approximations, which may limit accuracy in scenarios with complex posterior distributions. Additionally, they focus on lower-bound maximization which might sacrifice precision.
3 Expected-posterior estimation of the EIG gradient
@@
It follows a nested Monte Carlo estimator (30) given in Appendix A, using samples $\{({\mathb{y}}_{i},{\boldsymbol{\theta}}_{i})\}_{i=1:N}$ from the joint $p_{\boldsymbol{\xi}}$ and for each ${\mathb{y}}_{i}$, samples $\{{\boldsymbol{\theta}}^{\prime}_{i,j}\}_{j=1:M}$ from an MCMC procedure approximating the intractable posterior $p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$. Interestingly, expression (6) can also be recovered by setting the importance proposal $q({\boldsymbol{\theta}}^{\prime}|{\mathb{y}},{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ to $p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}},{\boldsymbol{\xi}})$ in (5), which provides a clear justification of why the choice of $q$ made in Goda et al., (2022) is relevant. Approaches by Goda et al., (2022) and Ao and Li, (2024) thus mainly differ in their choice of approximations for the posterior distribution. Using a Laplace approximation as in Goda et al., (2022) is relevant only if the posterior is unimodal, which may not be the case in practice. The MCMC version of Ao and Li, (2024) is then potentially more general but also more costly as it requires running $N$ times a MCMC sampler, targeting each time a different posterior $p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$. In the next paragraph, we introduce the expected posterior distribution and derive another, more computationally efficient, gradient expression.
Importance sampling EIG gradient estimator with an expected posterior proposal.
@@
We refer to $q_{{\boldsymbol{\xi}},N}$ as the expected posterior distribution, defined in a more general way in (13). It allows to assess the effect of a candidate design ${\boldsymbol{\xi}}$ on samples from the prior. It differs from a standard posterior as no real data ${\mathb{y}}$ obtained by running the experiment ${\boldsymbol{\xi}}$ is available during the optimization. We only have access to samples $\{({\boldsymbol{\theta}}_{i},{\mathb{y}}_{i})\}_{i=1:N}$ from the joint $p_{\boldsymbol{\xi}}$. The expected posterior can be seen as a distribution that takes into account all possible outcomes of a candidate experiment ${\boldsymbol{\xi}}$ given the samples $\{{\boldsymbol{\theta}}_{i}\}_{i=1:N}$ from the prior. This choice of $q_{{\boldsymbol{\xi}},N}$ is justified in Appendix B, using Lemma 2, proved therein. Lemma 2 shows that, for $\sum_{i=1}^{N}\!\nu_{i}=1$, $q_{{\boldsymbol{\xi}},N}$ is the distribution $q$ that minimizes the weighted sum of the KLs against each posterior $p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$, i.e. $\sum_{i=1}^{N}\!\nu_{i}\text{KL}(q,p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}}))$, leading to an efficient importance sampling proposal. It follows our new gradient estimator,
@@
Although with a reduced computational cost, computing gradients with (10) still requires an iterative sampling algorithm ideally run for a large number of iterations to reach satisfying approximations of the joint $p_{\boldsymbol{\xi}}$ and the expected posterior $q_{{\boldsymbol{\xi}},N}$. In static design, sampling from the joint is not generally difficult as the prior and the likelihood are assumed available but this becomes problematic in sequential design, as further detailed in Section 6.1. Sequential design is the setting to be kept in mind in this paper and in practice, the exact distributions are rarely reached. To assess the impact on gradient approximations, it is convenient to introduce, as in Marion et al., (2024), gradient operators. In the next section, we show how to adapt the formalism of Marion et al., (2024) to our BOED task.
@@ 4 EIG optimization through sampling
The intractability of the EIG and its gradient links optimization and sampling steps in a nested and costly way. Marion et al., (2024), in a different context, propose to adopt a bi-level optimization point of view to integrate sampling as an additional optimization step and derive simpler procedures.
@@
Denote $\Gamma$ a function from ${\cal P}({\boldsymbol{\Theta}}\times{\cal Y})\times{\cal P}({\boldsymbol{\Theta}})\times\mathbb{R}^{d}$ to $\mathbb{R}^{d}$, defined as,
@@
Expression (7) shows that $\nabla_{\boldsymbol{\xi}}I({\boldsymbol{\xi}})=\Gamma(p_{\boldsymbol{\xi}},q,{\boldsymbol{\xi}})$, where $q$ is a distribution $q({\boldsymbol{\theta}}^{\prime}|{\boldsymbol{\xi}})$ on ${\boldsymbol{\theta}}^{\prime}$ possibly depending on ${\boldsymbol{\xi}}$. The gradient estimator (10) corresponds then to $\nabla_{\boldsymbol{\xi}}I({\boldsymbol{\xi}})\approx\Gamma(\hat{p}_{\boldsymbol{\xi}},\hat{q}_{{\boldsymbol{\xi}},N},{\boldsymbol{\xi}})$, where $\hat{p}_{\boldsymbol{\xi}}=\sum_{i=1}^{N}\delta_{({\boldsymbol{\theta}}_{i},{\mathb{y}}_{i})}$ and $\hat{q}_{{\boldsymbol{\xi}},N}=\sum_{j=1}^{M}\delta_{{\boldsymbol{\theta}}^{\prime}_{j}}$. In general, sampling from $p_{\boldsymbol{\xi}}$, or its sequential counterpart, and $q_{{\boldsymbol{\xi}},N}$ is challenging and only possible through an iterative procedure. However, an interesting feature of our expected posterior is that it does not add additional sampling difficulties.
Expected posterior distribution.
@@
where $\rho$ is a measure on ${\cal Y}$. When $\rho({\mathb{y}})\!=\!\sum_{i=1}^{N}\!\nu_{i}\delta_{{\mathb{y}}_{i}}\!({\mathb{y}})$ with $\sum_{i=1}^{N}\!\nu_{i}\!=\!1$, we recover $q_{{\boldsymbol{\xi}},\rho}({\boldsymbol{\theta}})\!=\!q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}})$ in (9). The special structure of the expected posterior allows to sample from it using the same algorithmic structure to sample from a single posterior $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$. Indeed, the score of $q_{{\boldsymbol{\xi}},\rho}({\boldsymbol{\theta}})$ is linked to the posterior score,

### v2

@@ Abstract
Bayesian Optimal Experimental Design (BOED) is a powerful tool to reduce the cost of running a sequence of experiments. When based on the Expected Information Gain (EIG), design optimization corresponds to the maximization of some intractable expected contrast between prior and posterior distributions. Scaling this maximization to high dimensional and complex settings has been an issue due to BOED inherent computational complexity. In this work, we introduce a pooled posterior distribution with cost-effective sampling properties and provide a tractable access to the EIG contrast maximization via a new EIG gradient expression. Diffusion-based samplers are used to compute the dynamics of the pooled posterior and ideas from bi-level optimization are leveraged to derive an efficient joint sampling-optimization loop. The resulting efficiency gain allows to extend BOED to the well-tested generative capabilities of diffusion models. By incorporating generative models into the BOED framework, we expand its scope and its use in scenarios that were previously impractical. Numerical experiments and comparison with state-of-the-art methods show the potential of the approach.
@@ 1 Introduction
Designing optimal experiments can be critical in numerous applied contexts where experiments are constrained in terms of resources or more generally costly and limited. In this work, design is assumed to be characterized by some continuous parameters ${\boldsymbol{\xi}}\leavevmode\nobreak\ \in\leavevmode\nobreak\ {\cal E}\subset\mathbb{R}^{d}$, which refers to the experimental part, such as the choice of a measurement location, that can be controlled to optimize the experimental outcome. We consider a Bayesian setting in which the parameters of interest is ${\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}\subset\mathbb{R}^{m}$ and design is optimized to maximize the information gain on ${\boldsymbol{\theta}}$. Bayesian optimal experimental design (BOED) is not a new topic in statistics, see e.g. Chaloner and Verdinelli, (1995); Sebastiani and Wynn, (2000); Amzal et al., (2006) but has recently gained new interest with the use of machine learning techniques, see Rainforth et al., (2024); Huan et al., (2024) for recent reviews. The most common approach consists of maximizing the so-called expected information gain (EIG), which is a mutual information criterion that accounts for information via the Shannon’s entropy. Let $p({\boldsymbol{\theta}})$ denote a prior probability distribution and $p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ a likelihood defining the observation ${\mathb{y}}\in{\cal Y}$ generating process. The prior is assumed to be independent on ${\boldsymbol{\xi}}$ and $p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ available in closed-form. To our knowledge, all previous BOED approaches also assume that the prior is available in closed-form, a setting that we refer to as density-based BOED. In this work, by making BOED more computationally efficient, we open the first access to diffusion-based generative models and introduce data-based BOED when the prior is only available through samples. This broadens the scope of problems that can be tackled to a wide range of inverse problems (Daras et al.,, 2024).
The EIG, denoted below by $I$, admits several equivalent expressions, see e.g. Foster et al., (2019). It can be written as the expected loss in entropy when accounting for an observation ${\mathb{y}}$ at ${\boldsymbol{\xi}}$ (eq. (1)) or as a mutual information (MI) or expected Kullback-Leibler (KL) divergence (eq. (2)). Denoting $p_{\boldsymbol{\xi}}({\boldsymbol{\theta}},{\mathb{y}})=p({\boldsymbol{\theta}},{\mathb{y}}|{\boldsymbol{\xi}})$ the joint distribution of $({\boldsymbol{\theta}},{\mathb{Y}})$ and using $p({\boldsymbol{\theta}},{\mathb{y}}|{\boldsymbol{\xi}})=p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})p({\mathb{y}}|{\boldsymbol{\xi}})=p({\mathb{y}}|{\boldsymbol{\theta}},{\boldsymbol{\xi}})p({\boldsymbol{\theta}})$, it comes,
@@
In this work, we propose a novel 1-stage approach which leverages a sampling-as-optimization setting (Korba and Salim,, 2022; Marion et al.,, 2025) where sampling is seen as an optimization task over the space of probability distributions. We introduce a new EIG gradient expression (Section 3), which highlights the EIG gradient as a function of both the design and some sampling outcome. This fits into a bi-level optimization framework adapted to BOED in Section 4. So doing, at each step, both an estimation of the optimal design and samples from the current posterior distribution can be provided in a single loop described in Section 5. It results an efficient procedure that can handle both traditional density-based samplers and data-based samplers such as provided by the highly successful diffusion-based generative models. The resulting efficiency gain enables BOED applications at significantly larger scales than previously feasible, including inpainting problems ranging from computer vision to protein engineering and MRI (Quan et al.,, 2024; Yang et al.,, 2019; Aali et al.,, 2023). Figure 1 is an illustration on a $28\times 28$ image ${\boldsymbol{\theta}}$ reconstruction problem from $7\times 7$ sub-images centered at locations ${\boldsymbol{\xi}}$ to be selected, details in Section 6. For simpler notation, we first present our approach in the static design case. Adaptation to the sequential case is specified in Section 6 and all numerical examples are in the sequential setting.
@@ 2 Related work
We focus on gradient-based BOED for continuous problems. Applying a first-order method to solve (3) requires computing gradients of the EIG $I$, which are no more tractable than $I$ itself. Gradient-based BOED is generally based on stochastic gradient-type algorithms (see Section 4.3.2. in Huan et al., (2024)). This requires in principle unbiased gradient estimators, although stochastic approximation solutions using biased oracles have also been investigated, see e.g. Demidovich et al., (2023); Liu and Tajbakhsh, (2024). To meet this requirement, most stochastic gradient-based approaches start from an EIG lower bound that yields tractable unbiased gradient estimators. More specifically, EIG lower bounds have usually the advantage to remove the nested expectation issue, see e.g. Foster et al., (2019). In contrast, very few approaches focus on direct EIG gradient estimators. To our knowledge, this is only the case in Goda et al., (2022) and Ao and Li, (2024). Goda et al., (2022) propose an unbiased estimator of the EIG gradient using a randomized version of a multilevel nested Monte Carlo (MLMC) estimator from Rhee and Glynn, (2015). A different estimator is proposed by Ao and Li, (2024), who use MCMC samplers leading to biased estimators, for which the authors show empirically that the bias could be made negligible. In this work, we first show, in Section 3, that their two apparently different solutions actually only differ in the way the intractable posterior distribution is approximated. We then propose a third way to compute EIG gradients that is more computationally efficient and scales better to larger data volumes and sequential design contexts. This new expression makes use of a distribution that we introduce and name the pooled posterior distribution. This latter distribution has interesting sampling features that allow us to leverage score-based sampling techniques and connect to the so-called implicit diffusion framework of Marion et al., (2025). Our single loop procedure in Section 5 is inspired by Marion et al., (2025) and other recent developments in bi-level optimization (Yang et al.,, 2021; Dagréou et al.,, 2022; Hong et al.,, 2023). However, these latter settings do not cover doubly intractable objectives such as the EIG, which requires both appropriate gradient estimators and sampling operators, see our Sections 3 and 4. In BOED, efficient single loop procedures have been proposed by Foster et al., (2020) but they rely heavily on variational approximations, which may limit accuracy in scenarios with complex posterior distributions.
3 pooled-posterior estimation of the EIG gradient
@@
It follows a nested Monte Carlo estimator (30) given in Appendix A, using samples $\{({\mathb{y}}_{i},{\boldsymbol{\theta}}_{i})\}_{i=1:N}$ from the joint $p_{\boldsymbol{\xi}}$ and for each ${\mathb{y}}_{i}$, samples $\{{\boldsymbol{\theta}}^{\prime}_{i,j}\}_{j=1:M}$ from an MCMC procedure approximating the intractable posterior $p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$. Interestingly, expression (6) can also be recovered by setting the importance proposal $q({\boldsymbol{\theta}}^{\prime}|{\mathb{y}},{\boldsymbol{\theta}},{\boldsymbol{\xi}})$ to $p({\boldsymbol{\theta}}^{\prime}|{\mathb{y}},{\boldsymbol{\xi}})$ in (5), which provides a clear justification of why the choice of $q$ made in Goda et al., (2022) is relevant. Approaches by Goda et al., (2022) and Ao and Li, (2024) thus mainly differ in their choice of approximations for the posterior distribution. Using a Laplace approximation as in Goda et al., (2022) is relevant only if the posterior is unimodal, which may not be the case in practice. The MCMC version of Ao and Li, (2024) is then potentially more general but also more costly as it requires running $N$ times a MCMC sampler, targeting each time a different posterior $p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$. In the next paragraph, we introduce the pooled posterior distribution and derive another, more computationally efficient, gradient expression.
Importance sampling EIG gradient estimator with a pooled posterior proposal.
@@
We refer to $q_{{\boldsymbol{\xi}},N}$ as the pooled posterior distribution, defined in a more general way in (13). It allows to assess the effect of a candidate design ${\boldsymbol{\xi}}$ on samples from the prior. It differs from a standard posterior as no real data ${\mathb{y}}$ obtained by running the experiment ${\boldsymbol{\xi}}$ is available during the optimization. We only have access to samples $\{({\boldsymbol{\theta}}_{i},{\mathb{y}}_{i})\}_{i=1:N}$ from the joint $p_{\boldsymbol{\xi}}$. The pooled posterior can be seen as a distribution that takes into account all possible outcomes of a candidate experiment ${\boldsymbol{\xi}}$ given the samples $\{{\boldsymbol{\theta}}_{i}\}_{i=1:N}$ from the prior. This choice of $q_{{\boldsymbol{\xi}},N}$ is justified in Appendix B, using Lemma 2, proved therein. Lemma 2 shows that, for $\sum_{i=1}^{N}\!\nu_{i}=1$, $q_{{\boldsymbol{\xi}},N}$ is the distribution $q$ that minimizes the weighted sum of the KLs against each posterior $p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$, i.e. $\sum_{i=1}^{N}\!\nu_{i}\text{KL}(q,p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}}))$, leading to an efficient importance sampling proposal. It follows our new gradient estimator,
@@
Although with a reduced computational cost, computing gradients with (10) still requires an iterative sampling algorithm ideally run for a large number of iterations to reach satisfying approximations of the joint $p_{\boldsymbol{\xi}}$ and the pooled posterior $q_{{\boldsymbol{\xi}},N}$. In static design, sampling from the joint is not generally difficult as the prior and the likelihood are assumed available but this becomes problematic in sequential design, as further detailed in Section 6.1. Sequential design is the setting to be kept in mind in this paper and in practice, the exact distributions are rarely reached. To assess the impact on gradient approximations, it is convenient to introduce, as in Marion et al., (2025), gradient operators. In the next section, we show how to adapt the formalism of Marion et al., (2025) to our BOED task.
@@ 4 EIG optimization through sampling
To maximize the EIG using its gradient estimator (10), samples are needed from both the joint distribution $p_{\boldsymbol{\xi}}$ and the pooled posterior proposal $q_{{\boldsymbol{\xi}},N}$. If handled naively, it results a computationally expensive nested sampling-optimization loop where new samples from both distributions need to be generated at every update of the design parameter $\xi$. To derive more efficient procedures, we propose to adapt to BOED the framework of Marion et al., (2025) that integrates sampling and optimization into a single bi-level optimization loop. To do so, the EIG gradient $\nabla_{\xi}I(\xi)$ has first to be expressed as a function $\Gamma$ of three key components: the joint distribution $p_{\xi}$, the proposal distribution $q$, and the design parameter $\xi$ itself. Our choice of the pooled posterior as proposal distribution $q$ is then justified for its interesting sampling properties and the concept of sampling operator of Marion et al., (2025) is generalized to efficiently generate the samples needed to estimate our EIG gradient via $\Gamma$.
@@
Denote by $\Gamma$ a function from ${\cal P}({\boldsymbol{\Theta}}\times{\cal Y})\times{\cal P}({\boldsymbol{\Theta}})\times\mathbb{R}^{d}$ to $\mathbb{R}^{d}$, defined as,
@@
Expression (7) shows that $\nabla_{\boldsymbol{\xi}}I({\boldsymbol{\xi}})=\Gamma(p_{\boldsymbol{\xi}},q,{\boldsymbol{\xi}})$, where $q$ is a distribution $q({\boldsymbol{\theta}}^{\prime}|{\boldsymbol{\xi}})$ on ${\boldsymbol{\theta}}^{\prime}$ possibly depending on ${\boldsymbol{\xi}}$. The gradient estimator (10) corresponds then to $\nabla_{\boldsymbol{\xi}}I({\boldsymbol{\xi}})\approx\Gamma(\hat{p}_{\boldsymbol{\xi}},\hat{q}_{{\boldsymbol{\xi}},N},{\boldsymbol{\xi}})$, where $\hat{p}_{\boldsymbol{\xi}}=\sum_{i=1}^{N}\delta_{({\boldsymbol{\theta}}_{i},{\mathb{y}}_{i})}$ and $\hat{q}_{{\boldsymbol{\xi}},N}=\sum_{j=1}^{M}\delta_{{\boldsymbol{\theta}}^{\prime}_{j}}$. In general, sampling from $p_{\boldsymbol{\xi}}$, or its sequential counterpart, and $q_{{\boldsymbol{\xi}},N}$ is challenging and only possible through an iterative procedure. However, an interesting feature of our pooled posterior is that it does not add additional sampling difficulties.
Pooled posterior distribution.
@@
where $\rho$ is a measure on ${\cal Y}$. When $\rho({\mathb{y}})\!=\!\sum_{i=1}^{N}\!\nu_{i}\delta_{{\mathb{y}}_{i}}\!({\mathb{y}})$ with $\sum_{i=1}^{N}\!\nu_{i}\!=\!1$, we recover $q_{{\boldsymbol{\xi}},\rho}({\boldsymbol{\theta}})\!=\!q_{{\boldsymbol{\xi}},N}({\boldsymbol{\theta}})$ in (9). The special structure of the pooled posterior allows to sample from it using the same algorithmic structure to sample from a single posterior $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$. Indeed, the score of $q_{{\boldsymbol{\xi}},\rho}({\boldsymbol{\theta}})$ is linked to the posterior score,